# VilBias: A Study of Bias Detection through Linguistic and Visual Cues , presenting Annotation Strategies, Evaluation, and Key Challenges

### v3

@@ 3.1 Problem Definition
    Evaluations
3.2 Data Compilation
We selected a range of reputable and widely recognized news organizations for data collection, guided by prior research on media bias and source credibility [5, 63]. Our chosen outlets, CNN, BBC, The New York Times, and Al Jazeera, cover a broad spectrum of geographic regions, editorial policies, and audience trust profiles. Appendix LABEL:tab:news_sources provides detailed information about these sources.
Data were collected between May 6, 2023, and September 6, 2024, capturing evolving patterns of disinformation under shifting global events. To ensure compliant and ethical data practices, we adhered to privacy regulations, complied with website terms of service, and followed established guidelines [64].
Preprocessing focused on enhancing dataset quality and relevance. We removed duplicate articles to eliminate redundancy and filtered out multimedia-based items (e.g., URLs ending in /video, /slide, or /gallery) to concentrate on text-rich content. Articles with fewer than twenty sentences were also discarded to ensure sufficient textual depth for meaningful analysis. Each retained article was paired with its first encountered image, providing a consistent multimodal reference point.
Some of the images and text samples in this study are sourced from our prior work in same project [15]. However, the labeling scheme used here differs, as it incorporates multiple LLMs for annotation, enhancing robustness and reliability.
3.3 Multiple Annotations
To efficiently annotate a large, complex dataset, we integrated automated annotators in two categories:
  1. 1.
    LLMs for textual annotations, and
  2. 2.
    VLMs for multimodal (text-image) labels.
The prompt templates used for text-based annotations are provided in Appendix B.
A key feature of our approach is a two-step voting mechanism that increases annotation reliability. First, each LLM was queried three times per item to address stochastic variation inherent in LLMs [65], and the annotation from each LLM was determined by majority vote. Next, using multiple LLMs (three, in our case), the final annotation was decided by a second majority vote across these models.
Formally, let $A_{m}$ be the set of annotations from model $m$ after three queries. The final annotation $A_{\text{final}}$ is:
$A_{\text{final}}=\arg\max_{a\in A}\sum_{m\in M}\mathbb{1}(A_{m}=a),$
where $M$ is the set of LLMs (an odd number), $A$ is the set of all possible annotations, and $A_{m}$ is determined as:
$A_{m}=\arg\max_{a\in A}\sum_{i=1}^{3}\mathbb{1}(A_{m}^{(i)}=a),$
with $A_{m}^{(i)}$ being the $i$-th query response from model $m$. The indicator function $\mathbb{1}(A_{m}^{(i)}=a)$ equals 1 if the annotation matches $a$, and 0 otherwise. This voting approach mitigates stochastic variability within and across models, ensuring robust final annotation decisions.
Human Validation
Despite automation, human oversight remains essential for validating model outputs and calibrating the annotation pipeline. We enlisted 12 reviewers, comprising three undergraduates, four master’s students, two Ph.D. holders, and three senior researchers across computer science, media studies, political science, and linguistics. Two subject-matter experts led this process, guided by bias identification protocols outlined in Appendix B. These guidelines ensured that reviewers evaluated both linguistic and visual biases as defined by the U.S. Equal Employment Opportunity Commission [66].
Reviewers assessed a stratified sample of 10,000 annotations, classifying them as Acceptable, Needs Improvement, or Incorrect, and provided written rationales where necessary. This feedback informed iterative improvements to model prompts and annotation strategies. Disagreements were resolved through consensus discussions. After refining our approach, the inter-annotator agreement (IAA), calculated using Cohen’s Kappa, between automated labels and expert judgments reached 0.72, indicating substantial alignment.
Final Dataset
The resulting dataset is organized to support bias analyses across both textual and multimodal dimensions. Each article is assigned a stable, non-identifying unique identifier derived from the first ten digits of the SHA256 hash of its headline. Core fields include:
  • outlet, headline, article text (snippet), image, URL
To capture image semantics, we use a VLM (BLIP [67]) to generate a textual image description. The final dataset contains approximately 41,412 articles, each paired with its primary image. Two key annotation fields—nlp label and nlp-image label—indicate whether the text or the combined text-image pair is likely to exhibit bias. This schema enables diverse downstream tasks, from bias detection to explanatory modeling.
Table 3 summarizes key statistics and observed patterns. Figure 3 provides sample articles and their associated images, illustrating the multimodal scope of this collection.
Table 3: Data Analysis Summary.
Metric Value
Number of Rows 40,945
Number of Columns 7
Field Names unique_id, headline, url, article_text, image, nlp_label, nlp-image_label
Top Outlets (by count) Financial Times (4895), USA TODAY (3320), CNN (3067), The Guardian (2833), Forbes (2482)
Unique Values in nlp_label 2
Distribution of nlp_label {‘Bias’: 28,281, UnBias’: 12,664}
Unique Values in nlp-image_label 2
Distribution of nlp-image_label {‘Bias’: 22,974, ‘UnBias: 17,971}
Unique Images 40,945
Annotation Quality
We assess the annotation quality to address the research question: How accurate are annotations generated by LLM-based systems? To evaluate this, we selected a representative sample of 1000 records from our labeled dataset. This sample size was chosen to provide a statistically significant but manageable dataset for analysis and represents a balanced distribution between instances labeled to be biased or unbiased. The selected records were independently labeled by our team, who had no prior knowledge of the existing annotations. To assess the quality of LLM-generated annotations, we calculated precision, recall, F1 score, and accuracy, metrics widely used for evaluating classification performance.
Table 4: Performance metrics based on a majority vote from the dataset and human annotated sample, evaluated using a sample of 1000 examples. Metrics include precision, recall, F1 score, and accuracy for both text and multimodal labels.
Metric Text Label (%) Multimodal Label (%)
Precision 85.0 80.0
Recall 88.0 83.0
F1 Score 86.5 81.5
Accuracy 87.0 84.0
The results presented in Table 4 showcase the strong performance of LLM-based annotations, with all metrics surpassing 80%. This underscores the effectiveness of LLMs in handling both text-based and multimodal annotation tasks. Interestingly, text labels outperformed multimodal labels, achieving higher precision (85.0% vs. 80.0%), recall (88.0% vs. 83.0%), F1 score (86.5% vs. 81.5%), and accuracy (87.0% vs. 84.0%). These results highlight that multimodal labeling poses additional challenges, likely due to the complexities of integrating textual and visual information. Nevertheless, multimodal annotations remain competitive, with an accuracy of 84.0%, demonstrating their potential to capture subtle biases and richer context across modalities. The small gap between precision and recall for both text and multimodal labels reflects balanced and consistent performance, further validating the reliability of these annotations. Overall, these findings underscore the promise of multimodal models while also emphasizing the need for continued research to improve their ability to effectively leverage interactions between text and visuals.
4 Experimental Setup
4.1 Setting
For this work, we conducted our experiments using computing setup equipped with NVIDIA A100 Tensor Core GPUs to leverage efficient parallel processing capabilities for training and inference. The software environment included Python 3.10 with essential deep learning libraries such as PyTorch 2.0, Transformers by Hugging Face for model implementation, and OpenCV for image preprocessing. The models were trained and evaluated on Ubuntu 22.04 LTS running CUDA 12.1 for optimal GPU acceleration.
4.2 LLMs as Annotators
For text-only tasks, we utilized open source models such as Llama-3-8B-Instruct, Llama-3.1-8B-Instruct [68], and Mistral-7B-Instruct-v0.3 [69]. For text-image tasks, our approach integrated LLaVa-NeXT [70], Phi-3-Vision-128k-Instruct [71], and MiniCPM-V-2_6 [72]. To handle potential conflicts, we ensured an odd number of LLMs were used as annotators, enabling effective resolution of ties. This means that when there are disagreements among the annotators, the majority vote determines the final annotation, ensuring a definitive outcome without ambiguity. Our prompts for annotation are given in Appendix B.
4.3 Baseline Models and Evaluation Metrics
To conduct a comprehensive performance evaluation, we categorized models into SLMs and LLMs. In our study, SLMs encompass transformer architectures such as BERT [73], RoBERTa [74], DistilBERT [75], GPT-2 [76], and BART [77]. We also included multimodal configurations such as XLNet paired with VGG-19, BERT with ResNet-34, and BERT integrated with CLIP.
LLMs in our framework are represented by text-only models including LLama3.2-1B (both instruct and base variants) [78], Mistral-v0.3-instruct 7B [69], and Qwen2.5-7B-Instruct [79]. Furthermore, we examined VLMs such as InstructBlipV [80], Phi-3-vision-128k-instruct [71], Pixtral-12B-2409 [81], LLaVA-1.6 [82], and Llama-3.2-11B-Vision-Instruct [78]. This selection spans from compact, task-specific designs to versatile, billion-parameter frameworks, encompassing both text-only and multimodal platforms.
For evaluation, we tested each model under zero-shot and few-shot (5-shot) scenarios , and we also fine-tuned them using our entire dataset. We utilized an 80/10/10 split for training, validation, and testing subsets, respectively, ensuring a balanced representation of classes in each. Additionally, we implemented 5-fold cross-validation to verify the models’ robustness and employed stratified sampling to maintain uniform class distribution across the folds. Our metrics for performance assessment included standard ML metrics, such as precision, recall, F1-score, accuracy, and some qualitative evaluation through human validation.
4.4 Hyperparameters
All models were implemented using PyTorch, leveraging specific configurations to optimize performance and resource utilization. Key configurations include:
  • Data Type: torch.bfloat16 was used to enable mixed-precision training, striking a balance between memory efficiency and numerical accuracy.
  • Efficient Attention Mechanism: flash attention 2 was employed (where applicable) to accelerate attention computations while reducing memory overhead, particularly for larger models.
  • Device Mapping: Models were mapped to cuda-enabled devices to utilize GPU acceleration for faster computations.
Additionally, we employed quantization techniques when fine-tuning LLMs and VLMs to reduce memory requirements and speed up inference, particularly for large-scale models. For SLMs, regular fine-tuning without quantization was applied, as these models have a smaller parameter space and are computationally less demanding.
Table 5: Key Parameters for Training, Response Generation, and Evaluation
Parameter Details
Generation Temperature: 0.2–0.4, Maximum token limit: 1024
Batch Sizes Prompt Evaluation: 100
BERT Training: 32 (15 epochs, learning rate: $2\times 10^{-5}$, weight decay: 0.01, early stopping patience: 5)
Fine-Tuning: 1 (3 epochs, learning rate: $2\times 10^{-4}$, gradient accumulation: 1, max gradient norm: 0.3)
Learning Rate Schedulers Constant with warmup ratio of 0.03
Precision Mixed precision enabled (bf16=True)
Loss Functions CrossEntropyLoss (default)
Focal Loss (alpha=1, gamma=2) for specific tasks
Evaluation Metrics Accuracy, Precision, Recall, F1 Score, Confusion Matrices (Batch Size: 32)
Quantization 4-bit precision, double quantization enabled, nf4 quantization type
Environment and Reproducibility CUDA-enabled devices, CPU fallback when necessary
Random seed set to 0
Parallel tokenization disabled (os.environ["TOKENIZERS_PARALLELISM"] = "false")
5 Results
5.1 Overall Performance Evaluation
In this experiment, we investigate the comparative classification effectiveness of SLMs versus LLMs and VLMs. Our study specifically examines model performance across different configurations and modalities. We evaluated the models in two primary settings:
  1. 1.
    Text-only classification: Using small and large language models.
  2. 2.
    Multimodal classification: Employing small and large vision-language models.
Model Config Precision Recall F1 Score Accuracy
Small Language Models (Text-Only)
BERT-base-uncased FT 87.10% 86.50% 86.80% 86.90%
DistilBERT FT 84.30% 83.40% 83.85% 85.00%
RoBERTa-base FT 89.20% 87.90% 88.54% 89.50%
GPT-2 FT 85.50% 85.00% 85.24% 85.10%
BART FT 85.80% 85.70% 85.75% 85.80%
Large Language Models (Text-Only)
LLama3.2-1B-Instruct 0-shot 72.50% 73.40% 72.95% 67.80%
5-shot 76.20% 74.00% 75.08% 71.10%
IFT 84.00% 82.30% 83.14% 82.50%
Mistral-v0.3 7B-Instruct 0-shot 68.00% 65.00% 66.47% 67.50%
5-shot 80.10% 78.90% 79.49% 77.80%
IFT 86.20% 83.40% 84.78% 80.90%
Small Vision-Language Models (Text + Image)
XLNet + VGG-19 FT 72.00% 68.20% 70.05% 77.10%
BERT + ResNet-34 FT 75.80% 71.50% 73.59% 79.40%
BERT + CLIP FT 81.30% 73.40% 77.15% 84.20%
DistilBERT + CLIP FT 68.50% 64.00% 66.18% 74.90%
RoBERTa-base + CLIP FT 84.50% 81.40% 82.90% 83.60%
Large Vision-Language Models (Text + Image)
Phi-3-Vision-128k-Instruct 0-shot 70.40% 66.00% 68.12% 69.80%
5-shot 73.20% 71.00% 72.08% 70.50%
IFT 76.80% 78.10% 77.44% 74.00%
LLaVA-v1.6-Mistral-7B 0-shot 62.50% 60.80% 61.64% 62.70%
5-shot 68.10% 67.00% 67.54% 65.20%
IFT 75.40% 74.60% 75.00% 76.10%
LLama3.2-11B-Vision-Instruct 0-shot 65.00% 68.90% 66.88% 68.40%
5-shot 73.40% 74.20% 73.80% 72.10%
IFT 80.20% 81.90% 81.04% 82.50%
Table 6: Performance of small and large language models (text-only) and vision-language models (text + image) across configurations: 0-shot, 5-shot, fine-tuning (FT), and instruction fine-tuning (IFT). Metrics are presented in percentages.
Our evaluation reveals several key insights:
Text-only models: Table 6 provides an overview of how both SLMs and LLMs perform on text classification tasks. Among the SLMs, RoBERTa-base emerged as the best-performing model after fine-tuning, achieving an F1 score of 88.54% and an accuracy of 89.50%. Other SLMs, such as BERT and DistilBERT, also achieved strong results, although their simpler architectures and fewer parameters led to slightly lower performance.
For LLMs, Llama3.2 demonstrated robust performance, closely followed by Mistral. Their results underline the importance of training methodology and data size: instruction fine-tuning (IFT) outperformed both 5-shot and 0-shot approaches, showing that larger data samples and more demonstrations improve LLM training outcomes.
Although LLMs rely on patterns in training data to predict outputs, their predictions are not inherently probabilistic in the sense of randomness [83, 65]. During tasks like text classification, both SLMs and LLMs typically select the most likely output (argmax of the predicted distribution), which makes their behavior deterministic given fixed inputs and parameters. However, in tasks like text generation, probabilistic decoding methods (e.g., sampling with temperature or top-k sampling) may be used to introduce variability. Achieving a 100% success rate remains unrealistic for both SLMs and LLMs due to factors like limited training data, inherent ambiguities in language, and the complexity of interactions among parameters. While SLMs excel in specialized tasks due to focused optimization, LLMs rely on diverse and extensive training data to fully leverage their broader capabilities. These factors contribute to the modest performance gap observed between SLMs and LLMs for certain tasks.
Overall, these findings underscore that no single model or training approach is universally optimal. Instead, performance depends on a combination of factors, including the model’s architecture, the size and diversity of the training data, and the specifics of the fine-tuning process.
Multimodal Models: Table 6 presents the comparative performance of SLMs and large VLMs for multimodal classification tasks. Among small models, the combination of BERT with CLIP achieved the highest F1 score of 77.15% and an accuracy of 84.20% under the fine-tuning (FT) configuration. This result highlights the effectiveness of combining strong text encoders with vision models for multimodal tasks.
Large models demonstrated even greater potential, particularly with IFT. LLama3.2-11B-Vision-Instruct emerged as the best-performing model in this category, achieving an F1 score of 81.04% and an accuracy of 82.50%. The improvements underscore the importance of IFT in aligning multimodal capabilities, as it consistently outperformed both 0-shot and 5-shot configurations across models.
The results also reflect performance variations across configurations—0-shot, 5-shot, fine-tuning (FT), and IFT. For instance, small vision-language models like XLNet combined with VGG-19 showed relatively modest results, with an F1 score of 70.05% and accuracy of 77.10%. In contrast, larger models such as Phi-3-Vision-128k-Instruct demonstrated substantial improvements with IFT, reaching an F1 score of 77.44% and an accuracy of 74.00%. Similarly, LLaVA-v1.6-Mistral-7B showed the greatest gain when transitioning from 0-shot to IFT, reinforcing the value of tailored fine-tuning approaches for optimizing performance.
These results highlight the trade-offs between smaller and larger models, where small models often perform better in efficiency and simplicity, while large models leverage their scale and advanced training methodologies to achieve superior performance.
5.2 Ablation Studies
To better understand which components and training configurations most significantly impact model performance, we conducted a series of ablation experiments on fine-tuned both text-only and multimodal classification tasks. Specifically, we focused on three main aspects: data size, fine-tuning techniques, and multimodal fusion strategies.
5.2.1 Ablation 1: Training Data Size
We varied the size of the training dataset for a representative LLM (LLama3.2-1B-Instruct) and a SLM (RoBERTa-base) to examine how data quantity influences performance. The results in Figure LABEL:fig:ablation1 show that LLama3.2-1B outperforms RoBERTa-base in both F1 score and accuracy across all training data fractions. Additionally, LLama3.2-1B maintains better stability as the training data is reduced, indicating its robustness in low-resource scenarios, whereas RoBERTa-base exhibits a steeper decline in performance with reduced data availability. This shows that LLMs are more stable, and instruction fine-tuning further enhances their performance and resilience to limited training data.
5.2.2 Ablation 2: Fine-Tuning Techniques (Parameter-Efficient Methods)
Fine-tuning LLMs such as LLama3.2-1B and multimodal models like LLama3.2-11B-vision-Instruct can be computationally expensive due to the need to update all parameters during training. To address this, parameter-efficient fine-tuning (PEFT) [84] methods like Low-Rank Adaptation (LoRA) and Prefix-Tuning aim to reduce computational overhead by updating only a small subset of model parameters.
Full fine-tuning (FT) updates all model parameters during training, requiring high computational and memory resources but generally achieving the best performance. LoRA introduces low-rank trainable layers to the frozen pretrained weights, allowing significant parameter savings while maintaining performance. Prefix-Tuning appends trainable prefix embeddings to the input representations, leaving the main model frozen and thus reducing computational demands.
For this experiment, we used LLama3.2-1B-Instruct for text-only tasks and LLama3.2-11B-vision-Instruct for multimodal tasks (text + image). The models were trained on 80% of the dataset, with 20% held out for evaluation. The evaluation metrics used are pecision, recall, F1 score, and accuracy. The goal was to evaluate if these parameter-efficient methods could achieve competitive performance compared to full fine-tuning.
Table 7: Performance Comparison of Fine-Tuning Techniques for LLaMA 3.2 Models.
Model Method F1 Score Accuracy Drop (%)
LLama3.2-1B (Text) Full FT 83.1% 82.5% -
LoRA 82.3% 81.8% 0.8%
Prefix-Tuning 80.7% 80.5% 2.4%
LLama3.2-11B (Text + Vision) Full FT 81.7% 83.0% -
LoRA 81.0% 82.0% 0.7%
Prefix-Tuning 80.2% 81.2% 1.5%
Table 7 shows that full fine-tuning delivers the best performance across both text-only and multimodal tasks but comes at the cost of higher memory and computational requirements. LoRA achieves competitive performance with only a 0.7-0.8% drop in F1 score and accuracy, making it a highly practical alternative for resource-constrained environments. Prefix-tuning shows slightly more degradation, with a 1.5-2.4% drop in F1 score and accuracy, but remains viable when computational resources are extremely limited. These results demonstrate that PEFT methods can achieve near-parity with full fine-tuning while significantly reducing the computational overhead, making them suitable for deployment on edge devices or in scenarios with limited training infrastructure.
5.3 Qualitative Analysis and Bias Identification
In this section, we analyze potential biases in model predictions by examining how these models reason about certain statements. We elicit the chain-of-thought from LLMs and VLMs to determine whether their reasoning (1) relies on stereotypes or over-generalizations, or (2) correctly identifies or dismisses biases where relevant. We use our fine-tuned Llama3.2 ($1$B and $11$B), Mistral-v0.3-7B, and Phi3-vision model for generating predictions.
5.3.1 Text-based Qualitative Analysis and Bias Identification
Table 8 summarizes typical examples of each model’s responses. For each input statement, the model predicts whether there is Likely or Unlikely bias, then provides a brief rationale. We focus on two main aspects:
  1. 1.
    Misclassified Bias: Cases where the model either overestimates or underestimates the presence of bias due to stereotype-based reasoning or a failure to recognize marginalized perspectives.
  2. 2.
    Correct Classification: Cases where the model appropriately recognizes or dismisses bias based on the content and context of the statement.
Model Predic Example and Reasoning Bias Observed
Llama3.2-1B Bias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
Reasoning (model): “Nurses rarely contribute to surgical innovations; doctors usually lead.”
Issue: Stereotype-based reasoning. The model assumes nurses cannot be innovators, reinforcing traditional roles and overlooking individual competence.
Llama3.2-1B Unbias Input: “The committee led by women executives proposed significant policy reforms.”
Reasoning (model): “Focus on reforms; no inherent bias.”
Issue: Fails to acknowledge or explore the possibility of gender-based biases in leadership perceptions; too dismissive.
Llama3.2-11B Unbias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
Reasoning (model): “There is no explicit bias. Nurses can contribute to innovative practices, and the statement acknowledges that.”
Correct Classification: No stereotype is implied; the model recognizes nurses as capable innovators.
Mistral-7B-v0.3 Bias Input: “The startup founded by women faced challenges securing venture capital funding.”
Reasoning (model): “Historical funding gaps for women-led startups indicate bias.”
Issue & Strength: Identifies systemic funding biases (correct), but may overgeneralize by implying all such startups experience the same bias.
Mistral-7B-v0.3 Unbias Input: “The rural development program focused on uplifting marginalized communities.”
Reasoning (model): “Development programs are generally well-intentioned; bias is unlikely.”
Issue: Overlooks potential systemic biases in implementation or coverage, trivializing complexities of marginalized communities.
Mistral-7B-v0.3 Unbias Input: “The community initiative led by indigenous elders improved local education outcomes.”
Reasoning (model): “Recognizes indigenous leadership positively without implying stereotypes. No bias is evident.”
Correct Classification: The model captures a constructive scenario and does not assume negative stereotypes about indigenous communities.
Table 8: Binary bias classification (Bias/Unbias) with examples of both misclassified and correctly classified statements.
Our observations in Table 8 reveal that some models (e.g., smaller Llama3.2 variants) occasionally rely on stereotypes (e.g., downplaying a nurse’s capability). Models often overlook subtle socio-cultural biases, especially when they assume that “no explicit mention of bias implies no bias”. In some cases, models correctly identify known systemic biases (e.g., women-led startups facing funding challenges) or offer neutral interpretations when appropriate.
We also observe instances where models accurately identify or dismiss bias. For example, Llama3.2-11B correctly recognizes that highlighting a nurse’s innovative contributions is not inherently biased, and Mistral-7B-v0.3 appropriately identifies systemic gender biases for women-led startups. These cases show promise that with careful fine-tuning and context-awareness prompts, models can be guided to produce balanced and accurate assessments regarding bias.
Overall, these qualitative findings help us refine the models’ training objectives, such as incorporating more nuanced real-world scenarios into the fine-tuning data and instructing models to critically assess context.
5.3.2 Multimodal Qualitative Analysis and Bias Identification

### v5

@@ Appendix B Annotation Guidelines
  4. 4.
  5. 5.
    Bias Detection: Identify any systematic biases, such as stereotypes or skewed representations in text or vision-text pairs, that might emerge from the model’s outputs. Annotators were specifically instructed to consider the following bias categories, such as linguistic, visual, gender, racial, cutlural , disbility and age related biases.
  6. 6.
Appendix C Prompts
Prompt for Text-based Annotation prompt = f""" Analyze the provided text for evidence of bias. Identify the presence or absence of rhetorical techniques that contribute to bias in the news content. Provide a structured response to indicate how these techniques are used to influence the narrative. Text: {text} Rhetorical Techniques to Identify: 1. Emotional Appeal: Is the language emotionally charged to evoke strong reactions favoring one side? 2. Exaggeration and Hyperbole: Are claims exaggerated or overstated to emphasize one perspective? 3. Bias and Subjectivity: Does the text reflect polarized language or framing that favors one viewpoint? 4. Repetition: Are specific words, phrases, or ideas repeated to reinforce a particular narrative? 5. Specific Word Choices: Are charged or loaded terms used to subtly or overtly indicate bias? 6. Selective Presentation: Are facts or perspectives cherry-picked to highlight one side while ignoring others? 7. Omission of Context: Are important details or broader contexts omitted to skew the narrative? 8. Appeals to Authority: Are certain experts or authorities cited selectively to bolster a specific viewpoint? 9. Framing: Is the issue presented in a way that minimizes alternatives or discredits opposing perspectives? 10. Logical Fallacies: Are flawed arguments, such as straw man, ad hominem, or false dilemmas, employed to strengthen one side? 11. Tone: Is the tone neutral, or does it reflect sarcasm, disdain, or excessive praise for one perspective? 12. Imbalance in Representation: Are opposing views underrepresented or presented in a less favorable manner? 13. Use of Statistics or Data: Are statistics manipulated or selectively presented to support one narrative? Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation concise and structured, without additional explanations. """
Prompt for Text and Image-based Annotation prompt = f""" Analyze the provided text and associated image for bias. Evaluate how the text and image interact to reinforce or contradict the narrative and identify techniques used to promote a specific perspective. Text: {text} Image Description: {image_description} Evaluation Criteria: 1. Emotional Appeal: Does the text or image evoke strong emotions favoring one side? 2. Exaggeration and Hyperbole: Are claims or visual elements exaggerated to dramatize one perspective? 3. Bias and Subjectivity: Does the combination of text and image indicate clear alignment with one viewpoint? 4. Repetition: Are keywords, themes, or visual motifs repeated to reinforce a particular stance? 5. Specific Word Choices or Visual Cues: Are emotionally charged words or symbolic imagery used to promote bias? 6. Selective Presentation: Are key facts highlighted while opposing information is omitted or understated? 7. Omission of Context: Is critical background information missing in the text or image captions? 8. Appeals to Authority: Are the text or visuals aligned with selectively chosen or questionable authorities? 9. Framing: Is the issue framed to promote one perspective and discredit alternatives (e.g., through cropping, captions)? 10. Logical Fallacies: Are there flawed arguments or visual manipulations (e.g., misleading captions, doctored images)? 11. Tone: Is the tone in the text or visual style neutral, critical, or excessively favorable? 12. Imbalance in Representation: Are opposing perspectives underrepresented or negatively depicted? 13. Use of Symbolism: Are symbols, colors, or motifs in the image used to manipulate perception? 14. Use of Statistics or Data: Are charts, numbers, or data selectively presented or exaggerated? Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices or Visual Cues: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Symbolism: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation precise and avoid unnecessary explanations. """
LLM-as-Judge Prompt (ViLBias metrics: Reasoning Accuracy & Faithfulness) You are an impartial judge. Score a model’s open-ended rationale for a news text+image item on two binary metrics: (A) Reasoning Accuracy — does the rationale, if taken as true, correctly support the ground-truth label? (B) Faithfulness — is the rationale grounded in the provided input (text and/or image) with no hallucinated facts. Inputs - QUESTION: “Does this content exhibit bias? Please explain your reasoning.” - TEXT: <<<{TEXT SNIPPET}>>> - IMAGE: (rendered alongside; if unavailable, write “none”) - GROUND_TRUTH_LABEL: {biased | not_biased} - MODEL_PRED_LABEL: {biased | not_biased} - MODEL_RATIONALE: <<<{FREE-TEXT RATIONALE}>>> Judging Rules 1) Reasoning Accuracy (0/1): Assign 1 if the rationale’s claims, taken at face value, logically entail or strongly support the GROUND_TRUTH_LABEL. Penalize if it argues for the opposite label, is purely label restatement with no reason, or uses irrelevant logic. 2) Faithfulness (0/1): Assign 1 only if the rationale’s cited cues are verifiably present in TEXT and/or IMAGE. Disallow unstated assumptions, world knowledge, or fabricated details. When possible, quote minimal text spans and/or describe visual evidence (objects/attributes/relations); if no image, judge solely on text. 3) Edge Cases: - Empty, generic, or circular rationales (“it is biased because it is biased”) $\rightarrow$ ReasoningAccuracy=0, Faithfulness=0. - Correct label but incorrect/unsupported reasons $\rightarrow$ ReasoningAccuracy=0 or Faithfulness=0 as appropriate. - If MODEL_PRED_LABEL differs from GROUND_TRUTH_LABEL, you still judge the rationale against the ground truth. Output (strict JSON)
  "reasoning_accuracy": 0 or 1,
  "faithfulness": 0 or 1,
  "justification": "2-3 sentences: why each score was assigned"
Grading Hints - Typical textual bias cues: loaded wording, one-sided framing, selective emphasis/omission. - Typical visual bias cues: image choice, cropping/staging, emotive imagery, text–image mismatch. - Keep judgments short, precise, and grounded only in the provided inputs.
Figure A.1: LLM-as-Judge prompt for evaluation
Table A.1: Key settings for training, generation, and evaluation.
Generation Temperature 0.2–0.4; max tokens 1024
Batch / LR BERT: 32, 15 epochs, $2\!\times\!10^{-5}$, wd 0.01, ES patience 5
FT (VLM/LLM): batch 1, 3 epochs, $2\!\times\!10^{-4}$, grad acc 1, max grad norm 0.3
Scheduler Constant LR with 0.03 warmup ratio
Precision Mixed precision (bf16=True)
Loss CrossEntropy; Focal ($\alpha{=}1,\ \gamma{=}2$) where noted
Eval Accuracy, Precision, Recall, F1; confusion matrices (batch 32)
Quant. 4-bit (nf4) with double quantization
Reproduc. CUDA devices (CPU fallback), seed 0; TOKENIZERS_PARALLELISM=false
Table A.2: Binary classification sample results for various VLMs.